OpenStudy (ronaldo7):

√y+5 + √y +13 = 4 (the radical is over y+5 and y+13) Can someone help me with this

OpenStudy (mortonsalt):

The first thing you should try to do is eliminate the radicals over the y's. There's usually two ways you can do this: multiplying it with the conjugate or squaring both sides.

OpenStudy (ronaldo7):

the radicals are over both y+5 and y +13 not only y

OpenStudy (mortonsalt):

Yes, I understand. You can still do it this way, though.

OpenStudy (ronaldo7):

You cant just square root y and leave out the 5 or the 13

OpenStudy (mortonsalt):

The most sensible thing to do is to square both sides like so: \[(\sqrt{y+5} + \sqrt{y+13})^2 = 4^2\]

OpenStudy (mortonsalt):

This is what I mean. You square both sides.

OpenStudy (ronaldo7):

Ok I see where you're going

OpenStudy (mortonsalt):

You'll get something like: \[18+2y+2(\sqrt{y+5}+\sqrt{y+13})=16\]

OpenStudy (mortonsalt):

It should be easy from this point onward.

OpenStudy (ronaldo7):

No im soory i dont get it youkind of just added more stuff right?

OpenStudy (ronaldo7):

What i did i subtracted the rad y +5 and then squared both sides to get rid of the radicals

OpenStudy (mortonsalt):

\[(\sqrt{y+5} +\sqrt{y+13})(\sqrt{y+5} +\sqrt{y+13})=16\] Yes? Multiply the left hand side using the "FOIL" method.

OpenStudy (ronaldo7):

Not following you....

OpenStudy (mortonsalt):

Okay. Let's take it slow. Your equation is \[\sqrt{y+5}+\sqrt{y+13}=4\] You have to square both sides. \[(\sqrt{y+5}+\sqrt{y+13})^2 = 4^2\]

OpenStudy (mortonsalt):

Do you understand this so far?

OpenStudy (ronaldo7):

Ok but cant you also subtract the y+5 and then square it wouldnt that be easier

OpenStudy (ronaldo7):

I see what you mean but isnt this faster

OpenStudy (mortonsalt):

If you subtract the y+5, you would have to subtract it from BOTH SIDES. Which means you would end up with the same equation anyway.

OpenStudy (ronaldo7):

The rad y+5 so it can cancel one side

OpenStudy (mortonsalt):

\[\sqrt{y+5}+\sqrt{y+13}-\sqrt{y+15} = 4-\sqrt{y+5}\]

OpenStudy (mortonsalt):

Is this what you mean?

OpenStudy (ronaldo7):

the rad y +5 is suppose to cancel on the left side you didnt subtract it

OpenStudy (ronaldo7):

Or sorry i meant the y +3

OpenStudy (ronaldo7):

not y+5 i dont know where i got that one from

OpenStudy (kropot72):

It can be solved by using the following logic: \[0\le \sqrt{Y+5}<4\] \[1\le \sqrt{Y+13}\le4\] Note that Y can be a negative integer.

OpenStudy (ronaldo7):

I ended up with 24 - 8√y+13 sorry kropot i dont want to use inequalities

OpenStudy (ronaldo7):

is that right so far

OpenStudy (mortonsalt):

I'm not really sure what you're trying to do... But this should be the solution: \[(\sqrt{y+5}+\sqrt{y+13})^2=4^2\] \[18+2y+2\sqrt{(y+5)(y+13)}\] \[2\sqrt{(y+5)(y+13)}=-2-2y\] \[4(y+5)(y+13) = (-2-2y)^2\] \[ 4y^+72y+260=(-2-2y)^2\] \[4y^2+72y+260=4y^2+8y+4\] 64y+256=0 64(y+4)=0 y=-4

OpenStudy (mortonsalt):

Sorry, the second step should be =16

OpenStudy (mortonsalt):

The fourth one should be 4y^2 + 72y

OpenStudy (ronaldo7):

Let me do it on my own one more time and see what i come up with

OpenStudy (mortonsalt):

Okay.

OpenStudy (ronaldo7):

I got y to equal 525

OpenStudy (mortonsalt):

If you plug that value into the left side of your equation, you wouldn't get the right hand side (which is 4).

OpenStudy (ronaldo7):

Let me show you my work

OpenStudy (ronaldo7):

and sorry this taking so long

OpenStudy (ronaldo7):

\[√y+15 + √y+13 = 4 (√y+5)^{2} = (4- √y+13)^{2} y+5 = 16 - 8√y +13 + y + 13 y+5 +29 - 8√y +13 y = 24 - 8√y +13 y = 576 -64 + y + 13\] y = 525

OpenStudy (ronaldo7):

I see the problemthe y's will cancel out and leave an empty solution

OpenStudy (mortonsalt):

This solution isn't valid either. Remember that if you do something on one side, you have to do it in the other side as well. So if you multiply one side with whatever value, you have to do it on the other side.

OpenStudy (mortonsalt):

The solution you've given doesn't really make any sense. At least, to me.

OpenStudy (ronaldo7):

Yeah i see that now

OpenStudy (mortonsalt):

Are you going to be okay now? :)

OpenStudy (ronaldo7):

Yeah ill just ask my teacher tomorrow thank you anyway